OpenStudy (anonymous):

log(base3)3+log(base3)x=5

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

since the logs on the left are of the same base, can you rewrite that left side as one log base 3?

OpenStudy (anonymous):

you'll use this property: log_b (M) + log_b (N) = log_b (M*N)

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Yes, im sorry my brain is a little jumbled

OpenStudy (anonymous):

sawright...:)

OpenStudy (anonymous):

log_b (M) + log_b (N) = log_b (M*N) log_3 (3) + log_3 (x) = log_3 ( ) what do you think should go in the parenthesiss

OpenStudy (anonymous):

The 5

OpenStudy (anonymous):

no... try again... forget about that right side for now... we're just working on the left.

OpenStudy (anonymous):

On no its the 3 and x so its 3x =)

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

so we have: log_3 (3x) = 5 correct?

OpenStudy (anonymous):

Yes =) So we would divide noth sides by log-3

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

do you know how to rewrite this logarithm expression in it's exponential form?: log_3 (3x) = 5

OpenStudy (anonymous):

Yes =) the base is -3 and the exponent is 5. So it would be -3^5=3x??

OpenStudy (anonymous):

yes... btw, the base is positive 3. the line you see before that is just an underscore to denote the base.

OpenStudy (anonymous):

so 3^5 = 3x

OpenStudy (anonymous):

you're almost done.. what do you need to do to get x by itself?

OpenStudy (anonymous):

Divide by 3 on both sides, and I got 81 =)

OpenStudy (anonymous):

yep... ;)

OpenStudy (anonymous):

Oh yay! =) Thanks a ton!

OpenStudy (anonymous):

your good at these!!! nice work swimgirly...